OpenStudy (anonymous):

The lesson in my intermediate algebra class is rational expressions, the problem I've having trouble on is y = 5 sqrt y I have y^2 = 25y^2 and I'm unsure of where to go from there.

OpenStudy (anonymous):

\(y= 5\sqrt{y}\), square both sides \(y^2=25y\) hence \(y^2-25y =0\) factor y out \(y(y-25)=0\) that gives you y =0 and y =25

OpenStudy (anonymous):

When is it necessary to factor? Not for this problem but generally for rational expressions.

OpenStudy (anonymous):

when all terms in expression has the common "thing", like above, both the first term and the second term has y in

OpenStudy (anonymous):

Okay, thank you.